OpenStudy (anonymous):

What volume of carbon monoxide gas at STP reacts to produce 759 g of iron

OpenStudy (jebonna):

First write the equation out, which if I'm right should be this: Fe2O3 (s) + 3CO (g) -> 2Fe (s) + 3CO2 (g) From this we know the moles of each element in the equation. Now we can work out the volume of carbon monoxide. First work out the moles of Fe by using this equation: mass/molar mass 759g/55.8g mol-1 = 13.6mol (3.sig.fig) Now we need to deduce the moles. So the 13.6mol we have is for TWO moles of Fe. However, there are THREE moles of CO, meaning we have to make the answer we have now equal three moles, not two. To do this, first we can divide the answer we got by 2 to get the amount for 1 mol of a substance: 13.6/2 = 6.8mol Now times this by 3 to get the answer for 3 mol of a substance: 6.8 x 3 = 20.4mol of CO Now we can convert this into volume of CO. to do this we have to use this equation: moles x 24000cm^3 So for volume of CO we do: 20.4mol x 24000cm^3 = 489600cm^3 If you need the answer in dm^3, then divide what we got by 1000. 489600/1000 = 489.6dm^3 I hope this helps! If you get stuck on anything feel free to ask :)